factorize
4a^2 -4(squareroot)3 a+3
1
Expert's answer
2012-06-19T10:39:16-0400
we use (x-y)^2=x^2-2xy+y^2

4a^2
-4*sqrt(3)*a+3=(2a)^2-2*sqrt(3)*2a+(sqrt(3))^2=(2a-sqrt(3))^2